OpenStudy (anonymous):

Find the exact solution over the interval [0,2pie) 4cos ^2 x 8cos x +4=0

OpenStudy (ranga):

Is there a plus or a minus between 4cos^2(x) and 8cos(x) ?

OpenStudy (anonymous):

a plus

OpenStudy (ranga):

4cos^2(x) + 8cos(x) + 4 = 0 Divide throughout by 4 cos^2(x) + 2cos(x) + 1 = 0 This is a quadratic equation in cos(x) cos^2(x) + cos(x) + cos(x) + 1 = 0 cos(x)(cos(x) + 1) + 1(cos(x) + 1) = 0 (cos(x) + 1)(cos(x) + 1) = 0 (cos(x) + 1)^2 = 0 cos(x) + 1 = 0 cos(x) = -1 solve for x in the interval [0, 2pi)

OpenStudy (ranga):

cos(x) is minus 1 when x = pi.

OpenStudy (anonymous):

thank you! now I understand